# MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability

### v1

@@ Abstract
Large Language Model (LLM) agents deployed in complex real-world scenarios typically operate as spatially distributed entities. However, this physical dispersion constrains agents to limited local perception and finite temporal horizons. We characterize this bottleneck as spatiotemporal partial observability. Given such fragmented awareness, distributed agents struggle to coordinate efficiently. To bridge this gap, we introduce MACRO-LLM, LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability. The architecture addresses spatiotemporal constraints via three modules: (1) the CoProposer mitigates temporal uncertainty by verifying candidate actions via predictive rollouts; (2) the Negotiator overcomes spatial myopia by resolving conflicts through mean-field statistical aggregation; and (3) the Introspector ensures continuous adaptation by analyzing historical experience to refine strategies via semantic gradient descent. Extensive evaluations on two complex long-horizon tasks, cooperative adaptive cruise control and pandemic control, demonstrate that our framework effectively mitigates spatiotemporal partial observability through spatial and temporal strategies, enabling robust coordination.
@@ 1 Introduction
Large Language Model (LLM) agents have demonstrated remarkable human-like capabilities in problem-solving Zhu et al. (2023); Cui et al. (2025). While these agents successfully automate tasks across domains such as healthcare Wang et al. (2025); Khasentino et al. (2025) and education Zhao et al. (2025); Shi et al. (2025), single agents often struggle with complexity and scalability in long-horizon scenarios. Consequently, inspired by the team collaboration in human societies, researchers are actively exploring systems where multiple LLM agents cooperate toward common objectives Li et al. (2023a); Wang et al. (2024). This collective intelligence amplifies individual capabilities, enabling the effective resolution of complex reasoning tasks through coordination and negotiation Zou et al. (2025).
@@
However, a critical challenge emerges when deploying these systems in complex real-world scenarios Chen et al. (2025). Unlike centralized environments, agents in the physical world operate as distributed entities, which inherently impose spatiotemporal partial observability. Specifically, physical dispersion constrains agents to limited local perception, while the LLM’s fixed context window restricts their historical temporal information. For example, in autonomous driving, vehicles are bound by communication limits and privacy constraints, forcing them to optimize traffic flow using only local data. Therefore, addressing the partial observability stemming from real-world distribution is pivotal for enhancing collaborative efficiency and effectiveness.
@@
Specifically, partial observability manifests along two dimensions: spatial and temporal, as illustrated in Fig. 1(a). Spatially, constrained by communication limits and environmental complexity, each agent perceives only a fraction of the global state, preventing it from inferring the full context of the collective task. Temporally, agents face inherent uncertainty regarding future dynamics. Due to the stochastic nature of the environment and the unpredictable behaviors of others, the long-term consequences of current actions are difficult to predict. This forward-looking ambiguity, compounded by finite historical context, makes it challenging to align short-term decisions with long-term global objectives. Consequently, agents operate with a patchwork of incomplete spatial information and limited temporal understanding, resulting in suboptimal, inconsistent, or even conflicting actions that significantly degrade collective efficiency and effectiveness.
@@
Existing LLM-based Multi-Agent Systems (MAS) aim to expand their spatial perception through inter-agent communication, as shown in Fig. 1 (b). This is typically implemented via hierarchical structures that aggregate information to a central node Yu et al. (2024); Estornell et al. (2025) or fully-connected graphs that construct a pseudo-global view Chan et al. (2023). However, these methods often incur severe communication bottlenecks at central nodes or prohibitive overheads, limiting their scalability in large-scale, dynamically evolving environments. From a temporal perspective, while equipping agents with memory modules helps mitigate limited historical observations Park et al. (2023); Zhang et al. (2024), the inherent unpredictability of future states remains a critical challenge. This uncertainty makes it difficult for agents to align immediate actions with long-term goals and capture cascading temporal effects. Prior to the rise of LLM-based MAS, Multi-Agent Reinforcement Learning (MARL) provides a traditional framework for addressing partial observability. Nonetheless, MARL algorithms typically suffer from poor generalization and excessive training costs Huh and Mohapatra (2023). As illustrated in Fig. 7 for pandemic control tasks, training MARL models is computationally expensive. Such inflexibility and high resource demands severely restrict the feasibility of MARL solutions for complex real-world applications.
In this paper, we propose MACRO-LLM, an LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability, for efficient collaboration that can be generalized across diverse domains. To address spatial constraints, MACRO-LLM employs a neighborhood negotiation mechanism that combines iteratively refined semantic strategies with statistical spatial features, enabling a global understanding and coordinated decisions. To enhance temporal reasoning, MACRO-LLM integrates long-term future planning and short-term verification, continually updating strategies based on historical negotiation data. Specifically, each agent consists of three core modules: (1) the CoProposer, which generates collaborative action proposals and validates their feasibility via predictive rollouts to simulate short-term future states; (2) the Negotiator, which resolves conflicting proposals by analyzing both semantic strategies and mean-field statistical estimations of unobservable agents; and (3) the Introspector, which refines the agent’s temporal strategies via semantic gradient descent after executing each time-step action. Our main contributions are summarized as follows:
@@
  • Insightfully. To the best of our knowledge, this is the first LLM-based MAS to tackle partial observability by explicitly decomposing reasoning into spatial and temporal dimensions. We couple semantic reasoning with mean-field statistical features via neighborhood negotiation, enabling robust long-horizon coordination.
  • Technically. We develop MACRO-LLM, a synergistic architecture comprising three modules: a CoProposer for future unobservability, a Negotiator for spatial partial observability, and an Introspector for historical partial observability. Together, they facilitate the continuous refinement of spatiotemporal strategies through localized interactions.
  • Experimentally. We conduct comprehensive experiments on the Cooperative Adaptive Cruise Control (CACC) and Pandemic Control (PC) tasks, comparing against MARL and LLM-based baselines. Empirical results demonstrate that MACRO-LLM significantly outperforms existing methods, exhibiting superior coordination efficiency, scalability, and robustness against environmental dynamics.
@@ 2 Related Works
Existing LLM-based MAS have evolved from general collaboration frameworks Li et al. (2023a); Wu et al. (2023) to structured systems with persistent memory and reasoning capabilities Hong et al. (2023); Wang et al. (2023). To address spatial partial observability, recent works adopt either hierarchical aggregation, which centralizes information flow Yu et al. (2024); Estornell et al. (2025), or belief-based inference, which deduces hidden states via reasoning chains Li et al. (2023b); Davidson et al. (2024). However, these approaches often rely on ideal communication bandwidth or static topologies, failing to account for the dynamic coupling of spatial dispersion and temporal uncertainty. MACRO-LLM fills this gap by enabling decentralized spatiotemporal reasoning without reliance on central nodes or global views.
@@ 3 MACRO-LLM Framework
MACRO-LLM establishes a decentralized negotiation framework for MAS operating under spatiotemporal partial observability. As illustrated in Fig. 2, the architecture of agent $n$ comprises three synergistic modules: (1) CoProposer, for collaborative proposal generation; (2) Negotiator, for conflict resolution and spatial strategy updates; and (3) Introspector, for temporal strategy refinement via semantic gradient descent.
@@ 3.1 System Overview
We model the MAS as a peer-to-peer graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of agents and $\mathcal{E}$ represents the communication links. For any agent $n\in\mathcal{V}$, its neighborhood is defined as $\mathcal{N}_{n}=\{m\in\mathcal{V}|(n,m)\in\mathcal{E}\}$, consisting of all directly communicable neighbors. Reflecting real-world constraints, the system operates without a central coordinator.
The primary goal is to enable decentralized collaboration that maximizes global performance under local constraints. At initialization, each agent $n$ is provided with a system prompt defining the task, action space, state dynamics, and reward criteria.
@@ 3.2 CoProposer
The CoProposer module generates high-quality action proposals by integrating long-term objectives with immediate spatial constraints. At time step $t$, agent $n$ constructs a collaborative proposal $P_{n}^{t}$, comprising actions for itself and its observable neighbors $\mathcal{N}_{n}$ for subsequent negotiation.
@@
To incorporate spatial context, agent $n$ infers potential cascading effects originating from unobservable agents. By analyzing its local observation $o_{n}^{t}$ and the network topology $\mathcal{G}$, the agent forecasts multi-agent dynamics to formulate a spatial strategy $\Pi_{n,t}^{\text{spatial}}=\text{LLM}(o_{n}^{t},\mathcal{G})$.
@@
Leveraging current spatial and temporal strategies, agent $n$ first generates an initial proposal $P_{n}^{t,0}=(o_{n}^{t},a_{n}^{t,0},\{a_{n\rightarrow m}^{t,0}\})$ by identifying the most advantageous action for itself and inferring cooperative actions for its neighbors $\mathcal{N}_{n}$. Due to future uncertainty, this proposal requires validation through rollout-simulated verification. Agent $n$ first predicts its next observation $o_{n}^{t+1,0}$} and assesses whether this predicted future state constitutes an improvement to evaluate the proposal’s effectiveness. Then, this process iterates between proposal generation and next-observation verification over $k$ time steps, constituting a rollout.
Rollout: In our framework, a rollout is a multi-step simulation from time $t$ to $t+k$ that projects agent $n$’s future states and actions under its current strategy. The sequence of observations, actions, and rewards across these $k$ steps forms a trajectory, denoted as
$\tau_{n}=\Big((o_{n}^{t},a_{n}^{t},R_{n}^{t}),\dots,(o_{n}^{t+k},a_{n}^{t+k},R_{n}^{t+k})\Big).$ (1)
@@
To reduce computational complexity over long horizons, reward signal beyond $t+1$ is simplified to a binary indicator ($r=1$ if constraints are met, else 0). If a proposal violates constraints during the rollout, agent $n$ revises its action to generate a new candidate. This cycle repeats until a valid proposal is found or the maximum attempt limit is reached, at which point the candidate with the highest $\mathcal{R}(\tau_{n})$ is selected and broadcast to neighbors. This phase ensures each agent’s initial proposal maintains temporal consistency and spatial awareness for effective collaboration. The pseudocode is provided in Appendix B.
@@ 3.3 Negotiator
In decentralized environments, agents often generate proposals concurrently. Consequently, an agent must broadcast its own intent while processing incoming proposals from neighbors. This concurrency introduces conflicting actions, necessitating a negotiation mechanism to resolve conflicts and coordinate actions.
@@
To extend agents’ perception beyond observable limits, we complement semantic reasoning with statistical state features. In complex topologies, exchanging raw observations is inefficient due to bandwidth constraints. To address this, we introduce a protocol based on mean-field approximation Yang et al. (2018) to transmit aggregated statistical features.
The core idea of mean-field theory approximates complex pairwise interactions between an individual and all other agents in the environment can be approximated by the interaction between the individual and an aggregated “mean field” effect Lasry and Lions (2007). Accordingly, we represent the neighborhood influence through weighted statistics. For an agent $n$, the state $s_{m}^{t}$ of a neighbor $m$ is decomposed into a local mean field and a fluctuation term:
@@
Here, weights $w_{n,m}>0$ are determined by domain-specific relational factors, such as the spatial distance between agents. During negotiation, agents transmit these statistical features instead of raw data to characterize unobservable contexts. Once agent $n$ receives statistical features $(\mu_{m},\sigma_{m}^{2},W_{m})$ from a neighbor $m$, it updates the aggregated features incorporating its own state $s_{n}^{t}$ (with weight $w_{n}$) using the weighted Welford algorithm Welford (1962). This enables computationally efficient updates in one operation, the update equation is provided in Appendix C.
Proposal Confidence Assessment.
To evaluate proposal reliability, agent $n$ computes the pairwise differences between its own proposal $P_{n}^{t}$ and received neighbor proposals $\{P_{m}^{t}\}_{m\in\mathcal{N}_{n}}$. Local consensus is achieved if all differences fall below a predefined threshold $\epsilon$. Note that due to overlapping communication ranges, global consensus is only achieved when all agents in the network have reached local consensus and the negotiation is terminated. If consensus is not met, the agent analyzes the root causes of discrepancies, updates its spatial strategy by incorporating the received mean-field states, and assigns confidence weights to each neighbor’s proposal.
Proposal Regeneration.
The agent generates a refined proposal by integrating its original proposal with the weighted neighbor proposals, using the assigned confidence weights. This refined proposal then undergoes rollout verification as detailed in Sec. 3.2.
@@
A complete iteration of the CoProposer (Sec. 3.2) and Negotiator (Sec. 3.3) constitutes a negotiation round. MACRO-LLM typically performs $r$ rounds to exchange statistical features and proposals. Upon completion, agent $n$ synthesizes the proposal evolution to make a final confidence-weighted decision.

### v2

@@ Abstract
Large Language Model (LLM) agents deployed in complex real-world scenarios increasingly operate as spatially distributed entities. However, this physical dispersion constrains agents to limited local perception and finite temporal horizons. We characterize this bottleneck as spatiotemporal partial observability, where spatial and temporal limitations are fundamentally coupled: resolving spatial conflicts requires temporal reasoning about neighbors’ future actions, while temporal planning requires spatial context beyond local perception. To bridge this gap, we introduce MACRO-LLM, LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability. The architecture interleaves spatial and temporal reasoning within each decision cycle via three interdependent modules: (1) the CoProposer mitigates temporal uncertainty by verifying candidate actions via predictive rollouts; (2) the Negotiator overcomes spatial myopia by resolving conflicts through mean-field statistical aggregation, grounded in the CoProposer’s rollout rewards; and (3) the Introspector closes the reasoning loop by analyzing environmental drift and attributing performance changes to refine strategies. Extensive evaluations on two complex long-horizon tasks, cooperative platoon planning and pandemic control, demonstrate that our framework enables robust coordination under spatiotemporal partial observability.
@@ 1 Introduction
Large Language Model (LLM) agents have demonstrated remarkable human-like capabilities in problem-solving Zhu et al. (2023); Cui et al. (2025). While these agents successfully automate tasks across domains such as healthcare Khasentino et al. (2025) and education Zhao et al. (2025); Shi et al. (2025), single agents often struggle with complexity and scalability in long-horizon scenarios. Inspired by the team collaboration in human societies, researchers are actively exploring systems where multiple LLM agents coordinate to achieve collective goals Li et al. (2023a); Wang et al. (2024). This collective intelligence amplifies individual capabilities, enabling the effective resolution of complex reasoning tasks through coordination and negotiation Zou et al. (2025). However, most existing frameworks assume shared global context or unrestricted communication among agents. As LLM agents increasingly operate in physical-world scenarios, coordination under limited local perception remains underexplored.
@@
In such settings, agents operate as distributed entities Chen et al. (2025), which inherently impose spatiotemporal partial observability. Specifically, physical dispersion constrains agents to limited local perception, while the LLM’s fixed context window restricts their historical temporal information. This challenge is widespread: in pandemic control, districts must coordinate containment policies with only local infection data; in traffic coordination, vehicles optimize flow under communication limits and privacy constraints. Therefore, addressing the partial observability stemming from real-world distribution is pivotal for enhancing collaborative efficiency and effectiveness.
@@
Specifically, partial observability manifests along two dimensions: spatial and temporal, as illustrated in Fig. 1(a). Spatially, constrained by communication limits and environmental complexity, each agent perceives only a fraction of the global information, preventing it from inferring the full context of the collective task. Temporally, the stochastic environment and unpredictable behaviors of others make long-term consequences difficult to predict, while finite historical context further hinders alignment of short-term decisions with long-term objectives. Together, these constraints result in suboptimal or even conflicting actions that degrade collective efficiency and effectiveness.
@@
Existing LLM-based Multi-Agent Systems (MAS) aim to expand their spatial perception through inter-agent communication, as shown in Fig. 1 (b). This is typically implemented via hierarchical structures that aggregate information to a central node Yu et al. (2024); Estornell et al. (2025) or fully-connected graphs that construct a pseudo-global view Chan et al. (2023). However, these methods often incur communication bottlenecks at central nodes, limiting their scalability in large-scale, dynamically evolving environments. Temporally, while memory modules help mitigate limited historical observations Park et al. (2023); Zhang et al. (2024), the inherent unpredictability of future states remains unaddressed. This uncertainty makes it difficult for agents to align immediate actions with long-term goals and capture cascading temporal effects. Prior to LLM-based MAS, Multi-Agent Reinforcement Learning (MARL) provides a traditional framework for partial observability but suffers from poor generalization and excessive training costs Huh and Mohapatra (2023), limiting its feasibility for complex real-world applications.
In this paper, we propose MACRO-LLM, an LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability, generalizable across diverse domains. A key insight is that spatial and temporal partial observability are coupled: resolving spatial conflicts requires predicting how neighbors will act in the future, while planning over time requires spatial context beyond local perception. Existing approaches address these dimensions in isolation. MACRO-LLM bridges this gap through a negotiation-centric architecture where spatial and temporal reasoning are interleaved within each decision cycle via three interdependent modules: the CoProposer, which generates collaborative proposals validated via predictive rollouts to mitigate temporal uncertainty; the Negotiator, which resolves conflicts through mean-field statistical aggregation of unobservable agents to overcome spatial myopia; and the Introspector, which refines strategies via scenario-aware adaptive reflection to ensure continuous adaptation. Our main contributions are summarized as follows:
Insightful.
@@
We explicitly decompose partial observability into spatial and temporal dimensions and address them jointly through a unified decentralized negotiation framework that couples semantic reasoning with mean-field statistical features, enabling robust long-horizon coordination without central coordinators or global views.
Technical.
We develop MACRO-LLM with three interdependent modules forming a closed reasoning loop: the Negotiator’s confidence assessment consumes the CoProposer’s rollout rewards, and the Introspector’s revision signals reshape the strategies for subsequent cycles. This interdependence constitutes the core contribution beyond the individual techniques each module builds upon.
Experimental.
We evaluate on two complementary domains, Cooperative Platoon Planning (CPP) and Pandemic Coordination (PC), covering continuous and discrete actions, linear and graph topologies, fine-grained and long-horizon planning. Results demonstrate MACRO-LLM’s superior coordination performance, scalability, and robustness across six LLM backbones.
@@ 2 Related Works
Existing LLM-based MAS have evolved from general collaboration frameworks Li et al. (2023a); Wu et al. (2023) to structured systems with persistent memory and reasoning capabilities Hong et al. (2023); Wang et al. (2023). To address spatial partial observability, recent works adopt either hierarchical aggregation, which centralizes information flow Yu et al. (2024); Estornell et al. (2025), or belief-based inference, which deduces hidden states via reasoning chains Li et al. (2023b); Davidson et al. (2024). However, these approaches often rely on ideal communication bandwidth or static topologies, failing to account for the dynamic coupling of spatial dispersion and temporal uncertainty. Additionally, swarm-based optimization methods Feng et al. (2024); Zhang et al. (2025) evolve agent configurations through offline iterative search, yielding static configurations that cannot adapt to runtime dynamics and lack coordination mechanisms for resolving conflicts under partial observability (see Appendix A). MACRO-LLM fills this gap by enabling decentralized spatiotemporal reasoning without reliance on central nodes or global views.
@@ 3 MACRO-LLM Framework
MACRO-LLM establishes a decentralized negotiation framework for MAS operating under spatiotemporal partial observability. As illustrated in Fig. 2, the architecture of agent $n$ comprises three synergistic modules: (1) CoProposer, for collaborative proposal generation; (2) Negotiator, for conflict resolution and spatial strategy updates; and (3) Introspector, for temporal strategy refinement via scenario-aware adaptive reflection.
@@ 3.1 System Overview
We model the MAS as a peer-to-peer graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes agents and $\mathcal{E}$ communication links. For agent $n\in\mathcal{V}$, its neighborhood is $\mathcal{N}_{n}=\{m\in\mathcal{V}|(n,m)\in\mathcal{E}\}$. The system operates without a central coordinator. At initialization, each agent receives a system prompt defining the task, action space, state dynamics, and reward criteria. Architecturally, spatial and temporal reasoning are not separable. Evaluating an action requires predicting neighbors’ responses (temporal), but those predictions depend on unobservable states estimable only via mean-field spatial aggregation. MACRO-LLM operationalizes this dependency by interleaving the CoProposer and Negotiator within each negotiation round, with the Introspector providing cross-cycle feedback.
@@ 3.2 CoProposer
The CoProposer generates collaborative proposals $P_{n}^{t}$ comprising actions for agent $n$ and its observable neighbors $\mathcal{N}_{n}$ by integrating long-term objectives with immediate spatial constraints.
@@
To incorporate spatial context, agent $n$ infers potential cascading effects originating from unobservable agents. By analyzing its local observation $o_{n}^{t}$ and the network topology $\mathcal{G}$, the agent forecasts multi-agent dynamics to formulate a spatial strategy $\Pi_{n,t}^{\text{spatial}}$.
@@
Leveraging current spatial and temporal strategies, agent $n$ first generates an initial proposal $P_{n}^{t,0}=(o_{n}^{t},a_{n}^{t,0},\{a_{n\rightarrow m}^{t,0}\})$ by identifying the most advantageous action for itself and inferring cooperative actions for its neighbors $\mathcal{N}_{n}$. Due to future uncertainty, this proposal requires validation through rollout-simulated verification, performed via LLM inference without simulator access. Agent $n$ predicts $\hat{o}_{n}^{t+1,0}$ by reasoning over domain dynamics, $\Pi_{n,t}^{\text{temp}}$, and $\Pi_{n,t}^{\text{spatial}}$, then assesses whether the predicted state constitutes an improvement. This process iterates over $k$ steps, constituting a rollout.
Rollout: In our framework, a rollout is a multi-step LLM-based projection from time $t$ to $t+k$ that infers agent $n$’s future states and actions under its current strategy. The sequence of predicted observations, actions, and rewards across these $k$ steps forms a trajectory, denoted as $\tau_{n}=\Big((o_{n}^{t},a_{n}^{t},R_{n}^{t}),\dots,(o_{n}^{t+k},a_{n}^{t+k},R_{n}^{t+k})\Big)$.
@@
To reduce computational complexity over long horizons, reward signals beyond $t+1$ are simplified to a binary indicator ($r=1$ if constraints are met, else 0). If a proposal violates constraints during the rollout, agent $n$ revises its proposed action until a valid proposal is found or the maximum attempt limit is reached. The candidate with the highest $\mathcal{R}(\tau_{n})$ is selected and broadcast to neighbors. The pseudocode is provided in Appendix B.
@@ 3.3 Negotiator
In decentralized environments, agents generate proposals concurrently, inevitably introducing conflicts that necessitate a negotiation mechanism.
@@
To extend agents’ perception beyond observable limits, we complement semantic reasoning with statistical state features. In complex topologies, exchanging raw observations is inefficient due to bandwidth constraints. To address this, we introduce a protocol based on mean-field approximation to transmit aggregated statistical features.
The core idea of mean-field theory approximates complex pairwise interactions between an individual and all other agents in the environment by the interaction between the individual and an aggregated “mean field” effect Lasry and Lions (2007). Accordingly, we represent the neighborhood influence through weighted statistics. For an agent $n$, the state $s_{m}^{t}$ of a neighbor $m$ is decomposed into a local mean field and a fluctuation term:
@@
Here, weights $w_{n,m}>0$ are determined by domain-specific relational factors, such as the spatial distance, relative speed between agents. During negotiation, agents transmit these statistical features instead of raw data to characterize unobservable contexts. Once agent $n$ receives statistical features $(\mu_{m},\sigma_{m}^{2},W_{m})$ from a neighbor $m$, it updates the aggregated features incorporating its own state $s_{n}^{t}$ (with weight $w_{n}$) using the weighted Welford algorithm Welford (1962). Since each neighbor already aggregates its own neighborhood’s states into the transmitted statistics, this enables agent $n$ to incorporate spatial insights without requiring access to the raw states of all agents. The update equation is provided in Appendix F.
Eq. 1 defines the estimated cumulative reward $\mathcal{R}(\tau_{n})$ for agent $n$’s rollout trajectory $\tau_{n}$. The trajectory $\tau_{n}$ depends on the neighborhood states $\mathbf{s}_{\mathcal{N}_{n}}^{t}=(s_{1}^{t},\ldots,s_{|\mathcal{N}_{n}|}^{t})$, which influence the agent’s spatial strategy $\Pi_{n,t}^{\text{spatial}}$ (Sec. 3.2). Eq. 1 can be rewritten as $\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})\;\triangleq\;\mathcal{R}(\tau_{n})$ to expose the neighborhood dependence. Therefore, the error is denoted as $\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n})\right|$ i.e., the gap between the rollout reward computed with full neighborhood information $\mathbf{s}_{\mathcal{N}_{n}}^{t}$ and with the mean-field approximation $\boldsymbol{\mu}_{n}=(\mu_{n},\ldots,\mu_{n})$.
To justify transmitting $\sigma^{2}_{n}$ as an approximation quality indicator, we invoke Eqs. (7)(8) and Appendix B of Yang et al. (2018). In their formulation, pairwise Q-function interactions are approximated by the interaction with a mean agent, with the approximation error controlled by the variance of deviations from the mean. We apply the principle to analyze the theoretical properties of our mean-field aggregation, yielding the error bound:
$\left|\mathcal{R}_{n}(s^{t}_{\mathcal{N}_{n}})-\mathcal{R}_{n}(\mu_{n})\right|\leq\frac{L_{2}}{2w_{\min}}\cdot\sigma^{2}_{n},$ (4)
where $L_{2}$ is a smoothness constant determined by the continuity of the underlying domain dynamics and the boundedness of the state-action spaces, which hold broadly in physically grounded environments with continuous dynamics and finite horizons (a detailed discussion is provided in Appendix C). $\sigma^{2}_{n}$ controls the approximation error bound. In many topologies, agent $n$ acts as a cut vertex connecting disjoint subgroups (e.g., predecessor and follower in the CPP linear topology). Since these subgroups lack direct communication, their state distributions naturally diverge, making a single $\mu_{n}$ a poor summary. We prove that computing separate statistics per subgroup yields a provably tighter bound:
Theorem 1 (Tighter Bound via Topology-Aware Partitioning).
With the general error bound detailed in Eq. 4, computing separate weighted statistics per subgroup yields a tighter bound $\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\leq\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}|\mathcal{N}_{n}^{(p)}|\cdot\sigma_{n}^{2,(p)}\leq\frac{L_{2}}{2\,w_{\min}}\cdot|\mathcal{N}_{n}|\cdot\sigma_{n}^{2}$.
The proof is provided in Appendix D. Both bounds are controlled by $\sigma_{n}^{2}$. Thus $\sigma_{n}^{2}$ serves a dual role: (1) error indicator: small $\sigma_{n}^{2}$ guarantees accurate mean-field approximation; (2) uncertainty signal: large $\sigma_{n}^{2}$ prompts conservative strategies. During negotiation, agents progressively align their strategies, reducing $\sigma_{n}^{2}$ and tightening both bounds over successive rounds.
Proposal Confidence Assessment and Regeneration.
To evaluate proposal reliability, agent $n$ computes the pairwise differences between its own proposal $P_{n}^{t}$ and received neighbor proposals $\{P_{m}^{t}\}_{m\in\mathcal{N}_{n}}$. Local consensus between agents $n$ and $m$ is defined as $|P_{n}^{(q)}-P_{m}^{(q)}|<\delta$, where $\delta=\epsilon\cdot(a_{\max}-a_{\min})$ is proportional to the action space range for continuous actions, and $\delta=0$ for discrete action spaces requiring exact agreement. If consensus is not met, the agent analyzes discrepancies, updates its spatial strategy by incorporating the received mean-field states, and integrates weighted neighbor proposals to regenerate a refined candidate, which then undergoes rollout verification as detailed in Sec. 3.2.
@@
A complete iteration of the CoProposer (Sec. 3.2) and Negotiator (Sec. 3.3) constitutes a negotiation round. The primary purpose of multi-round negotiation is information propagation. Each round extends agents’ effective perception by one additional communication hop, with $r$ rounds providing coverage of $r+1$ hops. Local consensus serves as an early stopping criterion. Upon completion, agent $n$ produces a final decision by weighting each neighbor’s proposal by the confidence scores from proposal evaluation, selecting the highest-confidence action (or weighted combination for continuous actions). We provide Theorem 2, which establishes that proposal quality is monotonically non-decreasing across rounds.
Theorem 2 (Properties of the Negotiation Loop).
Let $\mathcal{G}$ be a connected graph and $r$ be the maximum negotiation rounds. The negotiation loop satisfies: (1) Bounded Termination: The negotiation terminates in at most $r$ rounds, either via early stopping when local consensus is reached ($\|P_{n}-P_{m}\|<\delta,\ \forall m\in\mathcal{N}_{n}$) or via the maximum round limit. (2) Monotonic Non-Decrease of Proposal Quality: $\mathcal{R}(\tau_{n}^{(\ell+1)})\geq\mathcal{R}(\tau_{n}^{(\ell)}),\quad\forall\,\ell=0,1,\ldots,r-1$.
The proof can be found in Appendix E. Regardless of whether agents reach consensus, the negotiation process terminates within $r$ rounds, and proposal quality is guaranteed to be monotonically non-decreasing across rounds, exhibiting a convergent trend toward improved coordination.